Buy Now and Gerald's EZ Furniture are two publicly traded rental companies. They reported the following in their financial statements (in millions of dollars, except per-share amounts and stock prices):

2016 2015 2016 2015
Net income $ 128 $ 180 $ 135 $ 189
Total stockholders' equity 1,127 1,150 1,333 1,235
Earnings per share 1.66 2.35 2.41 3.15
Stock price when annual
results reported
31.82 30.37 34.72 36.13

Required:

  1. 1-a. Compute the 2016 ROE for each company. Express ROE as a percentage. TIP: Remember that the bottom of the ROE ratio uses the average stockholders’ equity.

  2. 1-b. Which company appears to generate greater returns on stockholders’ equity in 2016?

  3. 2-a. Compute the 2016 P/E ratio for each company.

  4. 2-b. Which company do investors appear to value more?
Answer #1
1(a) ) (in millions of dollars) (in millions of dollars)
Buy Now Gerald EZ Furniture
2016 2016
Net income (A) $      128.00 $     135.00
Average Shareholders Equity (B) $   1,138.50 $  1,284.00
Return on Equity  = (A)/(B) 11.24% 10.51%
1(b)  BuyNow is geneating greater ROE in 2016.
2(a)
Market price per Share (A) $        31.82 $       34.72
EPS(B) $          1.66 $         2.41
P/E = (A)/(B) 19.17 14.41
2(b) = Buy Now have more Value